solve by substitution
x+y=5 and x=y+7

Answer by jim_thompson5910(35256)

x+y=5

y+7+y=5

2y+7=5

2y = 5-7

2y = -2

y = -2/2

y = -1

Now use this to find x. Your answer will be of the form (x,y)